OpenStudy (anonymous):

From the top of cliff 200 m above sea level the angles of depression to two fishing boats are 6.7 degrees and 8.2 degrees respectively. How far apart are the boats?

OpenStudy (dtan5457):

Draw it out.

OpenStudy (dtan5457):

So solve for each length based on the degrees perspective tan 6.7=n/200 tan 8.2=n/200 subtract larger from smaller when you get both values

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

thank youuu